OpenStudy (countrygirl16):

10th grade algebra help 5 questions check my 3 answers to see if I'm correct?

OpenStudy (countrygirl16):

1. x – 3 > –9 My answer: x >12 2. –2x + 1 ≤ –11 My answer: x >6 3. 10 < –3x + 1 My answer x >5 4. 2(x + 5) > 8x – 8 My answer is: Not sure 5. –2(x – 3) ≥ 5 – (x + 3) My answer is: Not sure

OpenStudy (countrygirl16):

am i correct on the first 3? and how do i do the last 2? im confused

OpenStudy (anonymous):

the first one is -12 i think...the other two are correct i think

OpenStudy (anonymous):

\[x-3>-9\] \[x>-9+3\] \[x>-6\]

OpenStudy (anonymous):

\[10 < –3x + 1 \]subtract 1 from both sides \[9<-3x\] divide both sides by \(-3\) and change the sense of the inequality\[-3>x\]

OpenStudy (anonymous):

satellite73 can you come help me when your done

OpenStudy (anonymous):

my question is the one below this

OpenStudy (anonymous):

\[2(x + 5) > 8x – 8\] distributive on the left \[2x+10>8x-8\] subtract \(2x\) from both sides \[10>6x-8\] add 8 to both sides \[18>6x\] divide both sides by 6\[3>x\]

OpenStudy (texaschic101):

1. x - 3 > - 9 x > -9 + 3 x > - 6 2. - 2x + 1 <= - 11 -2x <= - 11 - 1 -2x <= -12 x >= -12/-2 (the sign flips when dividing by a negative) x >= 6 3. 10 < -3x + 1 10 - 1 < - 3x 9 < - 3x 9/-3 > x -3 > x 4. 2(x + 5) > 8x - 8 2x + 10 > 8x - 8 10 + 8 > 8x - 2x 18 > 6x 3 > x 5. -2(x - 3) >= 5 - (x + 3) -2x + 6 >= 5 -x - 3 6 - 5 + 3 >= -x + 2x 4 >= x